OpenStudy (anonymous):

Find the first derivative of ln(cos(theta^2))

OpenStudy (anonymous):

\[2\theta*\tan(\theta^2)\]

OpenStudy (anonymous):

d/dx (log x) = 1/x

OpenStudy (anonymous):

d/dx ( log (f(x)) = [1/f(x)]* f`(x)